OpenStudy (anonymous):

a mother is 3tines as old as her daughter 6 years ago the sum of their ages were 36 find the age of the daughter and the mom

OpenStudy (anonymous):

3x=y x+y-12=36

OpenStudy (anonymous):

4x=48 x=12 y=36

OpenStudy (anonymous):

let daughter's present age = x years then mother's present age = 3x years 6 years ago daughter's age was = (x - 6) years and mother's age was = (3x - 6) years ATQ (x - 6) + (3x - 6) = 36 4x - 12 = 36 4x = 36 + 12 x = 48/4 x = 12 so daughter's present age = 12 and mother's present age = 36

OpenStudy (anonymous):

Very comprehensive answer, Harkirat.

OpenStudy (anonymous):

@AnimalAin thanks for the appreciation...☺

OpenStudy (anonymous):

@Harkirat U explained it well